OpenStudy (pottersheep):

Please help me prove this trig identity.... (secx - cosecx)/(tanx - cotx) = (tanx - cotx) /(secx - cosecx)

myininaya (myininaya):

i would write both sides in terms of sin(x) and cos(x) and clear the compound fractions

myininaya (myininaya):

and then you might have to use some other identities

OpenStudy (anonymous):

break up the LHS into \[\frac{sec x}{tan x - cot x} + \frac{-csc x}{tan x - cot x} \]

OpenStudy (pottersheep):

Oh....I'll try it that way then but Im not sure if I can reach the answer still.... And how can I change them to cos and sin? they are not squared :(

OpenStudy (anonymous):

my advice would be to multiply by one using \[\frac{sinx}{sinx} \]

OpenStudy (anonymous):

he means turn sec x into \[\frac{1}{cos x} \] etc...

OpenStudy (anonymous):

for multiplying by one like i said above, dont use sin...use something that will reduce everything...

OpenStudy (pottersheep):

hmmm

OpenStudy (pottersheep):

i tried turning tan into sin/cos and cot into cis/sin and then I got all tans and cos's too

OpenStudy (anonymous):

try taking everything in terms of cos and sin and then simplifying as much as you can. Don't leave any sec, csc, tan or cot

OpenStudy (mertsj):